Write a proportion for each phrase. Then solve, When nescessary, round to the nearest hundredth.
Question:
420 ft2 painted in 36 min; f ft2 painted in 30 min

Found 2 solutions by stanbon, JimMarshall:
Answer by stanbon(75887)

420/36 = f/30
f=(30)(420/36)
f=350 sq.ft.
Cheers,
Stan H.

Answer by JimMarshall(82)

420/36 = f/30
f=(30)(420/36)
f=350 sq.ft.
Ans